# SE Arena: Benchmarking Software Engineering Chatbots with Iterative Interactions

### v1

@@ Abstract
Foundation models (FMs), particularly large language models (LLMs), have shown significant promise in various software engineering (SE) tasks, including code generation, debugging, and requirement refinement. Despite these advances, existing evaluation frameworks are insufficient for assessing model performance in iterative, context-rich workflows characteristic of SE activities. To address this limitation, we introduce SE Arena, an interactive platform designed to evaluate SE-focused chatbots. SE Arena provides a transparent, open-source leaderboard, supports multi-round conversational workflows, and enables end-to-end model comparisons. Moreover, SE Arena incorporates a new feature called RepoChat, which automatically injects repository-related context (e.g., issues, commits, pull requests) into the conversation, further aligning evaluations with real-world development processes. This paper outlines the design and capabilities of SE Arena, emphasizing its potential to advance the evaluation and practical application of foundation models in software engineering.
@@ III-B Multi-Round Conversation Panel
All interactions remain anonymous, with chatbot identities concealed throughout the session. Votes are deemed valid only if (1) chatbot identities remain undisclosed in the conversation and (2) the submitted queries are relevant to software engineering tasks. A semi-automated periodic review process verifies the collected conversations, filtering out malicious or irrelevant voting behavior to ensure data reliability.
@@ IV Future Works
  • Advanced Multi-Round Evaluation Metrics: SE Arena intends to incorporate metrics specifically tailored to multi-round interactions. Examples include round-wise user feedback analysis to assess how models adapt and improve over successive turns.
@@
  • Developing a Guardrail for SE-Related Prompts: SE Arena plans to implement a guardrail mechanism to ensure that the prompts submitted are relevant to SE tasks. Non-SE-related prompts would be discarded, and the corresponding votes invalidated, maintaining the platform’s focus on providing meaningful insights into FM performance within software engineering.
@@
  • Expanding FM Coverage and Supporting Infrastructure: SE Arena aims to broaden coverage to include domain-specific models and multimodal foundation models, along with infrastructure for more complex SE tasks like web browsing and API integration.

### v2

# SE Arena: An Interactive Platform for Evaluating Foundation Models in Software Engineering

@@ Abstract
Foundation models (FMs), particularly large language models (LLMs), have shown significant promise in various software engineering (SE) tasks, including code generation, debugging, and requirement refinement. Despite these advances, existing evaluation frameworks are insufficient for assessing model performance in iterative, context-rich workflows characteristic of SE activities. To address this limitation, we introduce SE Arena, an interactive platform designed to evaluate SE-focused chatbots. SE Arena provides a transparent, open-source leaderboard, supports multi-round conversational workflows, and enables end-to-end model comparisons. Moreover, SE Arena incorporates a new feature called RepoChat, which automatically injects repository-related context (e.g., issues, commits, pull requests) into the conversation, further aligning evaluations with real-world development processes. This paper outlines the design and capabilities of SE Arena, emphasizing its potential to advance the evaluation and practical application of FMs in software engineering.
@@ III-B Multi-Round Conversation Panel
All interactions remain anonymous, with chatbot identities concealed throughout the session. Votes are deemed valid only if the submitted queries are relevant to SE tasks. To enforce this, SE Arena implements a guardrail mechanism that uses GPT-4o-mini [^6] https://openai.com/index/gpt-4o-mini-advancing-cost-efficient-intelligence to automatically filter out non-SE-related prompts at submission time, ensuring that evaluations remain focused on meaningful SE scenarios.
@@ IV Future Works
  • Developing Multi-round-oriented Evaluation Metrics: SE Arena intends to incorporate metrics specifically tailored to multi-round interactions. Examples include round-wise user feedback analysis to assess how models adapt and improve over successive turns.
@@
  • Expanding FM Coverage and Supporting Infrastructure: SE Arena aims to broaden coverage to include domain-specific models and multimodal foundation models, along with infrastructure for more complex SE tasks, including web browsing and API integration.